Use the following information for questions 1-7. A corporation has 9,000,000 shares of stock outstanding at a price of $40 per share. Net Income is $28,000,000 and they just paid a dividend of $2 and the dividend is expected to grow by 5% per year forever (Therefore next year’s dividend will be 2*(1.05) = $2.10). The stock has a beta of .9, the current risk free rate is 4%, and the market risk premium is 6%. The corporation also has 300,000 bonds outstanding with a price of $1,100 per bond. The bond has a coupon rate of 8% with semiannual interest payments, a face value of $1,000, and 13 years to go until maturity. The company plans on adding debt until they reach their target debt ratio of 70%. They expect their before-tax cost of debt to be 9% and their cost of equity to be 14% under this new capital structure. The tax rate is 25%

1. What is the CAPM required return on the corporation’s stock?

a) 9.4%          b) 10.25%       c) 11.3%         d) 12.2%        

2. What is the expected return on the corporation’s stock?

a) 9.4%           b) 10.25%       c) 11.3%         d) 12.2%        

3. What is the yield to maturity on the company’s debt?

a) 6.2%          b) 6.5%            c) 6.8%            d) 7.1%

4. What percent of their current market value capital structure is made up of debt?

a) 33%           b) 48%             c) 58%             d) 69%

5. What is their WACC using their target capital structure and expected costs of debt and equity?

a) 7.7%          b) 8.9%            c) 9.4%            d) 10.2%

6. Given the new cost of debt, what should be the new price of the bond?

a) $925          b) $960            c) $1,025         d) $1,175

7. Given the new cost of equity, what should be the new price of the stock?

a) $23.33       b) $27.25         c) $33.50         d) $36.67

Use the following information for questions 8-10: A company has Net Income of $40,000,000 and paid $18,000,000 in Dividends. They have 9,000,000 shares of stock at a price of $40 per share.

8. What is their sustainable growth rate?

a) 4.1%            b) 4.8%            c) 5.25%          d) 6.1%

9. How large can their capital budget be without having to issue new equity if they want a capital structure that is 30% equity?

a) $62,500,000            b) $67,250,000            c) $73,333,333            d) $87,500,000

10. What is their cost of equity if flotation costs are $5 per share if their growth rate is 5%?

a) 9.2%            b) 9.9%            c) 10.25%        d) 11%
Answer #1

1).

The formula for calculating the CAPM required Return:-

Ri = Rp + B(Rm)

Ri = Expected return of investment

Rf =Risk-free rate of Return : = 4%

β= Beta of Stock: = 0.9

Rm= Market Risk Premium : = 6%

The CAPM required return on Corporation's Stock

Ri= 4 + 0.9(6)

Ri= 9.4%

So, Option A

2).

Calculating Expected Return on stock using Dividend Growth Model:-

Dividend Growth Model

Di P = Ke-9

Po= Current Share Price: = $40

D1= Dividend paid next year: = $ 2.1

Ke= Cost of Equity/ Expected return

g= growth rate:= 5%

40= 2.1/(Ke- 0.05)

Ke-0.05 = 2.1/40

Ke= 10.25%

So, Option B

3).

Calculating Yield to Maturity (YTM) on company's Debt

+5-WA

C= Coupon Interest Payment ( since coupon is paid semi annually) : = $ 1000*8%*6/12 : = $ 40

F= Face Value/ Par Value

P= Current Price

n= period for years to maturity ( since semi annually paid) : = 13 years * 2 : = 26

40 + 1000-1100 26 YTM = 1000+1100

YTM = (40 - 3.846153)/1050

= 3.44%

Since, it is calculated semi annually, Yearly rate will be 6.88% (3.44*2)

So, Option C

4).

Share Price= $ 40

No. of Common Stock = 9,000,000

Total Market value of shares= $ 360,000,000

Current Bond Price= $ 1100

No. of Bonds= 300,000

Total Market value of bonds= $ 330,000,000

Total Market Value of capital Structure = $ 360 million + $ 330 million

= $ 690 million

Percent of Current market value of capital structure which is made up of debt is :-

= ($ 330 million/ $ 690 million)*100

= 47.82 % or 48%

So, Option B

Note- Since as per guidelines, I'm only allowed to answer maximum 4 sub-parts, First 4 has been answered.

If u like my answer do Up-vote as it will be motivating. Thanks